Please Expain the steps to divide
a^2-49/a-7

Answer by stanbon(75887)

Please Expain the steps to divide
a^2-49/a-7
---
Factor the numerator:
a^2-49 = (a-7)(a+7)
----
Problem is now:
[(a+7)(a-7)]/[a-7]
---
Cancel the factor common to numerator and denominator to get:
---
= a+7
===========
Cheers,
Stan H.
===========